A researcher wants to test whether there is a relationship between education level and smoking status. So the researcher collects some data and produces the following Minitab output. Find the missing p-value.

-------------------------------------------------------------
Tabulated statistics: smokingstatus2, edlevel2
Rows: smokingstatus2 Columns: edlevel2
H N All
D ? 30 67
? ? ?
S 6 ? ?
? ? ?
All ? ? ?
Cell Contents: Count
Expected count
Chi-Sq = 2.250, DF = ?, P-Value = ?

answer to 4 decimals.
Answer #1

here degree of freedom =(row-1)*(Column-1) =(2-1)*(2-1) =1

for 1 df and given chi square value , from excel p value =chidist(2.25,1)= 0.1336